### v1

# ars548_ros: An ARS 548 RDI radar driver for ROS2

Abstract
The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long range radar sensors with new RF antenna arrays, which offer digital beam forming. This radar measures independently the distance, speed and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation radarDesc . Unfortunately, there were not any drivers available for Linux systems to make the user able to analyze the data acquired from this sensor to the best of our knowledge. In this paper, we present a driver that is able to interpret the data from the ARS 548 RDI sensor and produce data in Robot Operation System version 2 (ROS2). Thus, this data can be stored, represented and analyzed by using the powerful tools offered by ROS2. Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity. We focus on the configuration of the sensor and the use of our driver and advanced filtering and representation tools, offering a video tutorial for these purposes. Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor for baseline purposes is available, so that the user can check the correctness of our driver.
keywords:
ROS2, ARS 548 RDI, Radar Sensor, FMPC radar
journal: SoftwareX
Code metadata
Current code version
1.0
Permanent link to code/repository used for this code version
https://github.com/robotics-upo/ars548_ros
Permanent link to Reproducible Capsule
Legal Code License
BSD 3-Clause License
Code versioning system used
Git
Software code languages, tools, and services used
C++, C.
Compilation requirements, operating environments & dependencies
Linux
If available Link to developer documentation/manual
https://github.com/robotics-upo/ars548_ros/blob/master/Readme.md
Support email for questions
Software metadata
Current software version
1.0
Permanent link to executables of this version
Permanent link to Reproducible Capsule
Legal Software License
List one of the approved licenses
Computing platforms/Operating Systems
Linux
Installation requirements
ROS2 Humble
Support email for questions
dalejo@us.es
1 Motivation and significance
Autonomous navigation in reduced visibility scenarios, i.e. scenarios with the presence of fog, rain, dust or smoke, is still a great challenge for robots, as the prevalent sensors used for obstacle detection in robotics, i.e. Light Detection and Ranging sensors (LiDAR), RGB and RGB-D cameras, can significantly reduce their precision or even become useless when operating under these conditions booker_jfr_2007 ; li_icra_2021 . Also, sensors onboard a smartphone can be used in robotic application or for indoor positioning thanks to applications such as GetSensorData GUTIERREZ2022101186 , but the performance of most onboard sensors would also drop in reduced visibility scenarios. Therefore, different sensors should be employed to allow autonomous robots to be deployed in reduced visibility scenarios.
One alternative can be the use of thermal infrared cameras for odometry and localization khattak2020keyframe ; saputra2020deeptio . In this paper, we focus on RAdio Detection And Ranging sensors (radar), as aerosols do not affect them because their wavelength is larger than most aerosol particles. Thus, radar sensors are able to penetrate rain, dust and smoke. However, using a larger wavelength also has drawbacks, as the measurements obtained by a radar sensor have worse distance and angular accuracy and resolution when compared to LiDAR measurements Mielle2019 .
Localization and navigation in reduced-visibility conditions is an active research field in autonomous car navigation. The vast majority of reduced-visibility datasets include RADAR information oxford_radar ; seeing_through_fog ; hong_radar_slam_2020 . A common approach is to perform data fusion of information perceived from LiDAR and RADAR sensors to ensure that the objects are detected in all circumstances. Unfortunately, the rotating radar sensors used in these datasets (such as the Navtech Radar CTS350-X) are too expensive and heavy to be included in a cost-effective robotic solution.
In the past years, advances in Frequency-Modulated Continuous Wave (FMCW) radar, which are based on the Doppler-effect to detect surrounding objects, have allowed the development of new radar sensors for robotic applications. Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of Doppler effect. This velocity information can be used to improve the odometry estimation in harsh environments radar_odometry_quist_2016 , which is fundamental for the safe autonomous navigation of the robots. Still, their use for 3D navigation is challenging due to the low resolution and narrow vertical Field of View (FoV) of inexpensive models, such as the IWR6843 from Texas Instruments.
More recently, the ARS 548 RDI sensor from Continental (see Figure 1) has been developed, which use more advanced modulation techniques to better estimate the position and velocity of nearby objects. We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios, as it is a good compromise solution between inexpensive but feature lacking FMCW radar sensors and 2D rotating radars used for automotive purposes.
In this paper, we describe the architecture of our software solution, highlighting its interface to Robotic Operating System v2 ros2 (ROS2), which allows us to use the representation, logging and processing tools offered by its community. We illustrate the good performance of our driver by including data collected from real experiments with the data of a OS1-32 Lidar from Ouster as a baseline of the data. The driver is distributed in a free license in the hope that it will be helpful to the community.
Figure 1: Continental ARS 548 RDI high resolution RADAR sensor.
2 Software description
We developed our driver with two main objectives in mind. On the one hand, our driver generates standard ROS2 messages such as PointCloud2 that includes information of the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects (see Figure 2). By using standard messages, the user can take advantage of existing representation and processing tools available in ROS2. On the other hand, we are interested in providing the user with all the information generated by the radar without any processing. To this end, we have created the following custom ROS 2 messages: ObjectList, DetectionList and Setup (see Figure 3), which are a direct translation of the data structures send by the sensor as described in the ARS 548 reference manual ars548_doc .
Figure 2: Basic flow diagram of ARS 548 driver.
2.1 Software architecture
The main components of our driver are listed on Figure 3. There are two main ROS2 packages in it, with the following main functionalities:
  • 1.
    ARS548_driver. This folder contains the libraries, code and functions needed for the actual driver to work. In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS548_messages), which are provided to the user via ROS2 middleware. On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
  • 2.
    ARS548_messages. In this folder we define the custom messages that are sent by our driver. These messages follow the data structure of the messages generated by the radar, as defined in ars548_doc .
Figure 3: Driver Architecture Scheme.
All of the code of the driver is available at our repository on GitHub repo .
2.2 Software functionalities
This section highlights the main functionalities provided by ARS 548 driver.
2.2.1 Reception and translation of messages.
The proposed driver is able to establish a UDP connection with the sensor and translate the different information from the sensor to custom and standard ROS2 messages, making the necessary endian translations by means of templated C++ functions.
With this functionality we provide the user with all the information of ARS548 as emitted by the sensors with our custom messages. Additionally, we provide standard messages which are easier to visualize and process with standard ROS2 packages.
2.2.2 Visualization of the results.
By generating standard ROS2 messages such as PointCloud2 and PoseArray, the user can easily represent the objects in the standard ROS2 visualizer (RViz2 rviz ). To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 3). You can see an example of representation of the radar measurements in Figure 4.
Figure 4: Example of object measurements captured by the ARS 548 RDI sensor, displayed in RViz.
2.2.3 Object filtering
Finally, our ARS 548 driver package provides the user with a customizable filter that is able to interpret our custom messages and emit filtered point clouds depending on the extended information generated by the sensor. As an example, see Section 3.2 in which we used a filter to only represent moving vehicles on the environment using the speed estimation of the sensor.
3 Illustrative examples
The driver presented in this paper has been experimentally tested in field experiments with the robotic ARCO platform arco (see Figure 5). In particular, we describe here two example scenarios. First, a traffic monitoring scenario, in which the velocity estimation offered by the radar sensor can be of great use to distinguish the road traffic from the background. Second, a typical robotic scenario, in which the robot should be capable of both generating a map of the environment and localizing itself on it, in a problem that is usually referred as Simultaneous Localization and Mapping (SLAM).
Figure 5: Our ARCO robotic platform equipped with both an ARS 548 RDI sensor and an OS1-32 LiDAR sensor. The necessary equipment for the experiments is pointed out on the Figure.
3.1 Experimental setup
Our ARCO robotic platform has been equipped with an ARS-548 sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 5). We used an MSI GF62 7RE Laptop with 32GB RAM and Intel Core i7 processor connected to the ARS 548 RDI sensor and the OS1-32 LiDAR sensor via Ethernet, with independent connections. For each experiment, we generated one or more ROSBag files with all the information from both sensors. In particular, with regards to ARS 548 RDI, we saved both our custom topics and the standard ones.
3.2 Experiment 1. Traffic monitoring
In this experiment, our ARCO robotic platform remained static in the proximity of a road. The objective here is to use the ARS 548 RDI sensor to monitor the road traffic nearby, taking advantage of its medium-long range (up to 300m) and its capability of measuring the speed of the detected objects. We use the object filtering feature of our driver obtaining an object point cloud that contains vehicles moving faster than 10 m/s, which are assumed as road traffic (see Figure 6).
Figure 6: Snapshot of Experiment 1. (left) Radar detections on blue, objects faster than 10 km/h on big white dots and LiDAR detections on multicolor. (right) Geolocalized radar measures represented over a map from OpensStreetMap. Red lines represent cumulative radar detections of two different vehicles over the map of the environment.
3.3 Experiment 2. SLAM
Work in progress.
4 Impact
This driver is published with a free software license in the hope of minimizing development efforts of all users of the ARS 548 RDI sensor that would like to implement ROS2-based applications with it. By doing so, the users can easily setup the driver and acquire its data with ROS2 interfaces and then focus on processing the data as needed for their own experiments, significantly speeding-up their research and development process.
Our main goal is to use this device for inspection robotics application carried out in reduced visibility environments, mainly due to weather phenomena such as rain, fog and snow or emergency situations involving smoke. These applications may include offshore oil&gas extraction plants inspection, sewer inspection and the inspection of transmission and electric power lines rs16020355 . By using this sensor, we can ensure that the inspection operations can be safely performed in all circumstances.
Another application field is autonomous driving. In fact, this is the main are of application of ARS 548 RDI devices. The idea here is to incorporate this sensor in an autonomous vehicle to provide it with detailed information of surrounding objects including the type of object or vehicle, and its relative velocity and acceleration. With this information, the user can design collision detection and resolution algorithms that guarantee the collision-free operation of the vehicle in spite of harsh weather or poor illumination circumstances.
5 Conclusions
In this project we describe a ROS2 driver to acquire and process all the data sent by the ARS 548 RDI radar, show some experiments monitoring the traffic to test its processing speed and precision and give examples where the use of a radar is necessary.
Acknowledgements
Authors would like to thank Fernando Amodeo for his invaluable help configuring the VLAN connection to the ARS 548 RDI sensor.