### v2

# ars548_ros: An ARS 548 RDI radar driver for ROS

Abstract
The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long range radar sensors with new RF antenna arrays, which offer digital beam forming. This radar measures independently the distance, speed and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1]. Unfortunately, to the best of our knowledge, there are no open source drivers available for Linux systems to enable users to analyze the data acquired by the sensor. In this paper, we present a driver that can interpret the data from the ARS 548 RDI sensor and make it available over the Robot Operating System versions 1 and 2 (ROS and ROS2). Thus, these data can be stored, represented, and analyzed using the powerful tools offered by ROS. Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity. We focus on the configuration of the sensor and the use of our driver including its filtering and representation tools. Besides, we offer a video tutorial to help in its configuration process. Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available, so that the user can check the correctness of our driver [2].
keywords: ROS ; Linux driver ; Radar Sensor
Code metadata
Current code version 1.0.0
Permanent link to code/repository used for this code version https://github.com/robotics-upo/ars548_ros
Permanent link to Reproducible Capsule
Legal Code License BSD 3-Clause License
Code versioning system used Git
Software code languages, tools, and services used C++, C.
Compilation requirements, operating environments & dependencies Linux
If available Link to developer documentation/manual https://github.com/robotics-upo/ars548_ros/blob/master/Readme.md
Support email for questions dalejo@us.es
Software metadata
Current software version 1.0.0
Permanent link to executables of this version
Permanent link to Reproducible Capsule
Legal Software License List one of the approved licenses
Computing platforms/Operating Systems Linux
Installation requirements ROS2 Humble or ROS Noetic
Support email for questions dalejo@us.es
1 Motivation and significance
Autonomous navigation in reduced visibility scenarios, i.e. scenarios with the presence of fog, rain, dust or smoke, is still a great challenge for robots, as the prevalent sensors used for obstacle detection in robotics, i.e. Light Detection and Ranging sensors (LiDAR), RGB and RGB-D cameras, can significantly reduce their precision or become useless when operating under these conditions [3, 4]. Also, sensors onboard a smartphone can be used in robotic applications thanks to applications such as GetSensorData [5], but the performance of most onboard sensors would also drop in reduced visibility scenarios. Therefore, different sensors should be employed to allow autonomous robots to be deployed in reduced visibility scenarios.
One alternative can be the use of thermal infrared cameras for odometry and localization [6, 7]. In this paper, we focus on RAdio Detection And Ranging sensors (radar), as aerosols do not affect them because their wavelength is larger than most aerosol particles. Thus, radar sensors are able to penetrate rain, dust and smoke. However, using a larger wavelength also has drawbacks, as the measurements obtained by a radar sensor have worse distance and angular accuracy and resolution when compared to LiDAR measurements [8].
Localization and navigation in reduced-visibility conditions is an active research field in autonomous car navigation. In fact, the vast majority of reduced-visibility datasets include radar information [9, 10, 11]. A common approach is to perform data fusion [12] of information perceived from LiDAR and radar sensors to ensure that the objects are detected in all circumstances. Unfortunately, the rotating radar sensors used in these datasets (such as the Navtech Radar CTS350-X) are too expensive and heavy to be included in a cost-effective robotic solution.
In the past years, advances in Frequency-Modulated Continuous Wave (FMCW) radar, which are based on the Doppler-effect to detect surrounding objects, have allowed the development of new radar sensors for robotic applications. Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of Doppler effect. This velocity information can be used to improve the odometry estimation in harsh environments [13], which is fundamental for the safe autonomous navigation of robots. Still, their use for 3D navigation is challenging due to the low resolution and narrow vertical Field of View (FoV) of inexpensive models, such as the IWR6843 from Texas Instruments.
More recently, the ARS 548 RDI sensor from Continental (see Figure 1) has been developed. It uses more advanced modulation techniques to better estimate the position and velocity of nearby objects. We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios, as it is a good compromise solution between inexpensive but feature lacking FMCW radar sensors and 2D rotating radars used for automotive purposes. Unfortunately, we did not find any open source driver that would allow us to acquire and process the data acquired from this sensor.
Figure 1: Continental ARS 548 RDI high resolution radar sensor.
In this paper, we describe the architecture of the ars548_ros package. We highlight its interface to both versions of the Robotic Operating System [14] (ROS), which are becoming the standard for robotic development at least for academic purposes, even though ROS2 is more industrial focused than its predecessor, ROS. To develop ROS and ROS2 interfaces is very convenient as it allows us to use its representation, logging and processing tools. Hereinafter, we will use ROS when we refer to both versions of ROS. We illustrate the good performance of our driver by including data collected from real experiments with measurements of a OS1-32 LiDAR sensor from Ouster as a baseline. Finally, the driver is distributed in the BSD 3-Clause License in the hope that it will be helpful to the community.
2 Software description
We developed our driver with two main objectives in mind. On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information of the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects. By using standard messages, the user can take advantage of existing representation and processing tools available in ROS. On the other hand, we are interested in providing the user with all the information generated by the radar without any processing. To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures send by the sensor as described in the ARS 548 reference manual [15]. The messages have been included in Appendix A for the sake of completeness. Finally, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
Figure 2: a) Basic flow diagram of ARS 548 driver. b) Example application of the filter: detection of objects moving faster than a threshold velocity.
2.1 Software architecture
Our solution consists of two ROS packages, with the following main functionalities:
  • ARS548_driver. This folder contains the libraries, code and functions needed for the actual driver to work. In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS548_messages), which are provided to the user via ROS middleware. On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
  • ARS548_messages. In this folder we define the custom messages that are sent by our driver. These messages follow the data structure of the messages generated by the radar, as defined in [15].
The full driver source code is publicly available in our GitHub repository [16]. Also, please refer to Appendix B for the additional contents associated with the paper.
2.2 Software functionalities
This section highlights the main functionalities provided by ARS 548 driver.
2.2.1 Reception and translation of messages.
The proposed driver is able to establish a UDP connection with the sensor and translate the different information from the sensor to custom and standard ROS messages, making the necessary endian translations by means of templatized C++ functions.
With this functionality we provide the user with all the information from ARS548 as emitted by the sensors with our custom messages. Additionally, we provide standard messages which are easier to visualize and process with standard ROS packages.
2.2.2 Visualization of the results.
By generating the PointCloud2 and PoseArray standard ROS messages, the user can easily represent the objects in the standard ROS visualizer (RViz [17]). To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 2a). You can see an example of representation of the radar measurements in Figure 3.
Figure 3: Radar detection measurements captured by the ARS 548 RDI sensor in blue and OS1 liDAR measurements in red. Please note the alignment between LiDAR and radar measurements.
2.2.3 Object filtering
The ars548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor. The available information is defined in the Object message described in Appendix A. The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details. As an example, see Section 3.2 in which we used a filter to represent only moving vehicles in the environment using the speed estimation from the sensor.
3 Illustrative examples
The driver presented in this paper has been experimentally tested in field experiments with the robotic ARCO platform [18] (see Figure 4). In particular, we describe here two example scenarios. First, a traffic monitoring scenario, in which the velocity estimation offered by the radar sensor can be of great use to distinguish the road traffic from the background. Second, a typical robotic scenario, in which the robot should be capable of both generating a map of the environment and localizing itself on it, in a problem that is usually referred as Simultaneous Localization and Mapping (SLAM).
Figure 4: Our ARCO robotic platform equipped with both an ARS 548 RDI sensor and an OS1-32 LiDAR sensor. The necessary equipment for the experiments is pointed out on the Figure.
3.1 Experimental setup
Our ARCO robotic platform has been equipped with an ARS 548 RDI sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 4). We used an MSI GF62 7RE Laptop with 32GB RAM and Intel Core i7 processor connected to both sensors via Ethernet, with independent connections. For each experiment, we generated one or more ROSBag files with all the information from both sensors. In particular, with regards to ARS 548 RDI, we saved both our custom topics and the standard ones. The dataset is available at [2].
3.2 Experiment 1. Traffic monitoring
In this experiment, our ARCO robotic platform remained static in the proximity of a road. The objective here is to use the ARS 548 RDI sensor to monitor the road traffic nearby, taking advantage of its medium-long range (up to 300m) and its capability of measuring the speed of the detected objects. We use the object filtering feature of our driver (see Figure 2b) obtaining an object point cloud that contains vehicles moving faster than 10 km/h, which are assumed as road traffic (see Figure 5).
Figure 5: Snapshot of Experiment 1. (left) Radar detections on blue, Radar objects on red, Radar object moving faster than 10 km/h on a green sphere and LiDAR measurements on white. (right) Geolocalized radar measurements represented over a map from OpenStreetMap [19]. Red lines represent cumulative radar detections of two different vehicles over the map of the environment.
3.3 Experiment 2. SLAM
In the second experiment, our ARCO robotic platform navigates within an open environment while estimating its trajectory through the information received by the radar. Figure 3 illustrates the alignment between the measurements from our driver and the LiDAR, verifying its proper functioning.
To validate the results of this experiment, we have employed the odometry generated by the LeGO-LOAM package [20] as a baseline, which relies on the LiDAR sensor for data acquisition (see Figure 6). The data acquired from both the radar and LiDAR sensors will be used for future odometry and SLAM techniques, ensuring accurate mapping and localization, especially in hazardous scenarios where radar is employed.
Figure 6: The baseline trajectory generated by the LeGO-LOAM package is represented in large green dots over the generated map projection in 2D. The LiDAR measurements are represented in red dots, with small green dots representing features detected by Lego-LOAM.
4 Impact
This open source driver has been published under the BSD 3-Clause License in the hope of minimizing the development efforts of all users of the ARS 548 RDI sensor who would like to implement ROS-based applications with it. By doing so, the users can easily setup the driver and acquire its data with ROS interfaces and then focus on processing the data as needed for their own experiments, significantly speeding up their research and development process.
Our main goal is to use this device for inspection robotics application carried out in reduced visibility environments, mainly due to weather phenomena such as rain, fog and snow or emergency situations involving smoke. These applications may include the inspection of different facilities such as offshore oil&gas extraction plants, sewers and transmission and electric power lines [21]. By using this sensor, we can ensure that inspection operations can be performed safely under all circumstances.
Another application field is autonomous driving. In fact, this is the main area of application for the ARS 548 RDI devices. The idea here is to incorporate this sensor into an autonomous vehicle to provide it with detailed information of surrounding objects including the type of object or vehicle and its relative velocity and acceleration. With this information, the user can design collision detection and resolution algorithms that guarantee the collision-free operation of the vehicle in spite of harsh weather or poor illumination circumstances.
5 Conclusions
This paper described a ROS driver to configure and interface with the ARS 548 RDI radar sensor. The driver works on both ROS and ROS2 versions, maximizing its impact on the community. The driver has been tested in two different field experiments, showcasing the advanced information from the sensor and the filtering capabilities of our driver.
Future work will focus on developing new radar odometry approaches, and LiDAR-radar data fusion mechanisms.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
All the data used in this paper is openly available at [2].
Acknowledgements
Authors would like to thank Fernando Amodeo for his invaluable help. This work has been supported by the grants INSERTION PID2021-127648OB-C31 and RATEC PDC2022-133643-C21. Both grants have been funded by the Spanish Ministry of Science, Innovation and Universities with grant number 10.13039/501100011033 and “European Union NextGenerationEU/PRTR”.
Appendix A . Custom ARS 548 messages
Below you can find the definition of the custom ROS messages in ars548_messages.
Detection.msg
1float32 f_azimuthangle #Unaligned Detection Azimuth Angle
2float32 f_azimuthanglestd #Azimuth Angle Std
3uint8 u_invalidflags #Detection Invalid Flags
4float32 f_elevationangle #Unaligned Detection Elevation Angle
5float32 f_elevationanglestd #Elevation Angle Std
6float32 f_range #Detection Radial Distance
7float32 f_rangestd #Radial Distance Std
8float32 f_rangerate #Detection Radial Velocity
9float32 f_rangeratestd #Radial Velocity Std
10int8 s_rcs #Detection RCS
11uint16 u_measurementid #Detection ID
12uint8 u_positivepredictivevalue #Existence Probability
13uint8 u_classification #Detection Classification
14uint8 u_multitargetprobabilitym #Multi-Target Probability
15uint16 u_objectid #Associated Object ID
16uint8 u_ambiguityflag #Probability for resolved velocity ambiguity
17uint16 u_sortindex #tbd
DetectionList.msg
1std_msgs/Header header
2uint64 crc #Checksum (E2E Profile 7) (Reserved)
3uint32 length #Len (E2E Profile 7) (Reserved)
4uint32 sqc #SQC (E2E Profile 7) (Reserved)
5uint32 dataid #Data ID (E2E Profile 7) (Reserved)
6uint32 timestamp_nanoseconds #Timestamp Nanoseconds
7uint32 timestamp_seconds #Timestamp Seconds
8uint8 timestamp_syncstatus #Timestamp Sync Status
9uint32 eventdataqualifier #Event Data Qualifier (unused)
10uint8 extendedqualifier #Extended Qualifier (unused)
11uint16 origin_invalidflags #Sensor Position Invalid flags (unused)
12float32 origin_xpos #Sensor X Position with reference to rear axle
13float32 origin_xstd #Sensor X Position STD (unused)
14float32 origin_ypos #Sensor Y Position with reference to rear axle
15float32 origin_ystd #Sensor Y Position STD (unused)
16float32 origin_zpos #Sensor Z Position with reference to rear axle
17float32 origin_zstd #Sensor Z Position STD (unused)
18float32 origin_roll #Sensor Roll Angle (unused)
19float32 origin_rollstd #Sensor Roll Angle STD (unused)
20float32 origin_pitch #Sensor Pitch Angle with alignment correction
21float32 origin_pitchstd #Sensor Pitch Angle STD
22float32 origin_yaw #Sensor Yaw Angle with alignment correction
23float32 origin_yawstd #Sensor Yaw Angle STD
24uint8 list_invalidflags #Invalid flags (unused)
25Detection[800] list_detections #Detection Array
26float32 list_radveldomain_min #Ambiguity free Doppler velocity range Min
27float32 list_radveldomain_max #Ambiguity free Doppler velocity range Max
28uint32 list_numofdetections #Number of Detections
29float32 aln_azimuthcorrection #Azimuth Alignment Correction
30float32 aln_elevationcorrection #Elevation Alignment Correction
31uint8 aln_status #Status of alignment
Object.msg
1uint16 u_statussensor #tbd
2uint32 u_id #Unique ID of object
3uint16 u_age #Age of object
4uint8 u_statusmeasurement #Object Status
5uint8 u_statusmovement #Object Movement Status
6uint16 u_position_invalidflags #tbd
7uint8 u_position_reference #Reference point position
8float32 u_position_x #X Position of reference point
9float32 u_position_x_std #X Position Std
10float32 u_position_y #Y Position of reference point
11float32 u_position_y_std #Y Position Std
12float32 u_position_z #Z Position of reference point
13float32 u_position_z_std #Z Position Std
14float32 u_position_covariancexy #Covariance X Y
15float32 u_position_orientation #Object Orientation
16float32 u_position_orientation_std #Orientation Std
17uint8 u_existence_invalidflags #unused
18float32 u_existence_probability #Probability of Existence
19float32 u_existence_ppv #unused
20uint8 u_classification_car #Car Classification Probability
21uint8 u_classification_truck #Truck Classification Probability
22uint8 u_classification_motorcycle #Motorcycle Classification Probability
23uint8 u_classification_bicycle #Bicycle Classification Probability
24uint8 u_classification_pedestrian #Pedestrian Classification Probability
25uint8 u_classification_animal #Animal Classification Probability
26uint8 u_classification_hazard #Hazard Classification Probability
27uint8 u_classification_unknown #Unknown Classification Probability
28uint8 u_classification_overdrivable #unused
29uint8 u_classification_underdrivable #unused
30uint8 u_dynamics_absvel_invalidflags #unused
31float32 f_dynamics_absvel_x #X Absolute Velocity
32float32 f_dynamics_absvel_x_std #X Absolute Velocity Std
33float32 f_dynamics_absvel_y #Y Absolute Velocity
34float32 f_dynamics_absvel_y_std #Y Absolute Velocity Std
35float32 f_dynamics_absvel_covariancexy #Covariance Absolute Velocity X Y
36uint8 u_dynamics_relvel_invalidflags #unused
37float32 f_dynamics_relvel_x #X Relative Velocity
38float32 f_dynamics_relvel_x_std #X Relative Velocity Std
39float32 f_dynamics_relvel_y #Y Relative Velocity
40float32 f_dynamics_relvel_y_std #Y Relative Velocity Std
41float32 f_dynamics_relvel_covariancexy #Covariance Relative Velocity X Y
42uint8 u_dynamics_absaccel_invalidflags #unused
43float32 f_dynamics_absaccel_x #X Absolute Acceleration
44float32 f_dynamics_absaccel_x_std #X Absolute Acceleration Std
45float32 f_dynamics_absaccel_y #Y Absolute Acceleration
46float32 f_dynamics_absaccel_y_std #Y Absolute Acceleration Std
47float32 f_dynamics_absaccel_covariancexy #Covariance Absolute Acceleration X Y
48uint8 u_dynamics_relaccel_invalidflags #unused
49float32 f_dynamics_relaccel_x #X Relative Acceleration
50float32 f_dynamics_relaccel_x_std #X Relative Acceleration Std
51float32 f_dynamics_relaccel_y #Y Relative Acceleration
52float32 f_dynamics_relaccel_y_std #Y Relative Acceleration Std
53float32 f_dynamics_relaccel_covariancexy #Covariance Relative Acceleration X Y
54uint8 u_dynamics_orientation_invalidflags #unused
55float32 u_dynamics_orientation_rate_mean #Object Orientation Rate
56float32 u_dynamics_orientation_rate_std #Orientation Rate Std
57uint32 u_shape_length_status #(unused) Shape Length Status
58uint8 u_shape_length_edge_invalidflags #(unused) Invalid Flags Shape Length
59float32 u_shape_length_edge_mean #Mean Shape Length
60float32 u_shape_length_edge_std #(unused) Shape Length Std
61uint32 u_shape_width_status #(unused) Shape Width Status
62uint8 u_shape_width_edge_invalidflags #(unused) Invalid Flags Shape Width
63float32 u_shape_width_edge_mean #Mean Shape Width
64float32 u_shape_width_edge_std #(unused) Shape Width Std
ObjectList.msg
1std_msgs/Header header
2uint64 crc #Checksum (E2E Profile 7) (Reserved)
3uint32 length #Len (E2E Profile 7) (Reserved)
4uint32 sqc #SQC (E2E Profile 7) (Reserved)
5uint32 dataid #Data ID (E2E Profile 7) (Reserved)
6uint32 timestamp_nanoseconds #Timestamp Nanoseconds
7uint32 timestamp_seconds #Timestamp Seconds
8uint8 timestamp_syncstatus #Timestamp Sync Status
9uint32 eventdataqualifier #(unused) Event Data Qualifier
10uint8 extendedqualifier #(unused) Extended Qualifier
11uint8 objectlist_numofobjects #Number of Objects
12Object[50] objectlist_objects #Object Array
Status.msg
1uint32 timestamp_nanoseconds #Timestamp Nanoseconds
2uint32 timestamp_seconds #Timestamp Seconds
3uint8 timestamp_syncstatus #Timestamp Sync Status
4uint8 swversion_major #Software version (major)
5uint8 swversion_minor #Software version (minor)
6uint8 swversion_patch #Software version (patch)
7float32 longitudinal #Longitudinal sensor position (AUTOSAR)
8float32 lateral #Lateral sensor position (AUTOSAR)
9float32 vertical #Vertical sensor position (AUTOSAR)
10float32 yaw #Sensor yaw angle (AUTOSAR)
11float32 pitch #Sensor pitch angle (AUTOSAR)
12uint8 plugorientation #Orientation of plug
13float32 length #Vehicle length
14float32 width #Vehicle width
15float32 height #Vehicle height
16float32 wheelbase #Vehicle wheelbase
17uint16 maximundistance #Maximum detection distance
18uint8 frequencyslot #Center frequency
19uint8 cycletime #Cycle time
20uint8 timeslot #Cycle offset
21uint8 hcc #Country code
22uint8 powersave_standstill #Power saving in standstill
23uint32 sensoripaddress_0 #Sensor IP address
24uint32 sensoripaddress_1 #Reserved
25uint8 configurationcounter #Counter that counts up if new configuration has been received and accepted
26uint8 status_longitudinalvelocity #Signals if current VDY is OK or timed out
27uint8 status_longitudinalacceleration #Signals if current VDY is OK or timed out
28uint8 status_lateralacceleration #Signals if current VDY is OK or timed out
29uint8 status_yawrate #Signals if current VDY is OK or timed out
30uint8 status_steeringangle #Signals if current VDY is OK or timed out
31uint8 status_drivingdirection #Signals if current VDY is OK or timed out
32uint8 status_characteristicspeed #(Unused) Signals if current VDY is OK or timed out.
33uint8 status_radarstatus #Signals if Radar Status is OK
34uint8 status_voltagestatus # Bitfield to report under- and overvoltage errors
35uint8 status_temperaturestatus #Bitfield to report under- and overtemperature errors
36uint8 status_blockagestatus #Current blockage state and blockage self test state.
Appendix B . Supplementary data
The video tutorial for configuring the sensor and the driver can be found at [22]. The repository of the driver can be found at [16] and the acquired dataset can be found at [2].